# Understanding and Mitigating Membership Inference Risks of Neural Ordinary Differential Equations

### v1

@@ Appendix B Differential Privacy of Gaussian Mechanism
The following result is proved in (Dwork and Roth, 2014). Since it is crucial to the results in the body, a slightly different proof is provided here for completeness.
Proposition B.1 (Theorem A.1 in (Dwork and Roth, 2014)).
@@
$\Pr$ $\left[~{}Z>t-\frac{\varepsilon}{2t}\right]$
$\leq\frac{1}{\sqrt{2\pi}}\left(t-\frac{\varepsilon}{2t}\right)^{-1}\exp\left(-\frac{1}{2}\left(t-\frac{\varepsilon}{2t}\right)^{2}\right)<\frac{\delta}{2},$
@@
Note that, by enforcing $t\geq 1$ and $10<\varepsilon<$ 1, the first term on the left-hand side is nonnegative when
$t-\frac{\varepsilon}{2t}\geq t-\frac{1}{2}\geq 1,$
Therefore, it suffices to choose $t\geq 3/2$ to ensure positivity. On the other hand, the second term is increasing in $t$ when $t>\sqrt{\varepsilon/2}$ and decreasing in $\varepsilon$ when $\varepsilon<2t^{2}$, so it follows that we may ensure
$\left(t-\frac{\varepsilon}{2t}\right)^{2}$ $=t^{2}-\varepsilon+\left(\frac{\varepsilon}{2t}\right)^{2}\geq t^{2}-1+1\cdot 3^{-2}$
$=t^{2}-\frac{8}{9}>2\ln\left(\frac{1}{\delta}\sqrt{\frac{2}{\pi}}\right).$
@@
$t^{2}$ $>2\ln\left(\frac{1}{\delta}\sqrt{\frac{2}{\pi}}\right)+\ln\left(e^{\frac{8}{9}}\right)$
$=2\ln\left(\frac{1}{\delta}\right)+\ln\left(e^{\frac{8}{9}}\frac{2}{\pi}\right),$
and since $(2/\pi)e^{8/9}<1.25^{2}$, this condition is satisfied when
$t^{2}>2\ln\left(\frac{1.25}{\delta}\right).$
@@
This shows that the privacy loss random variable is bounded above by $\varepsilon$ with probability $1-\delta$ on ${}q,~{}q^{\prime}$ whenever $\sigma\geq\sqrt{2\ln\left(1.25/\delta\right)}\allowbreak(\left|~{}v\right|/\varepsilon)$. To remove the dependence on ${}q,~{}q^{\prime}$, it suffices to choose $\sigma\geq\sqrt{2\ln\left(1.25/\delta\right)}(S_{~{}f}/\varepsilon)$. ∎
Appendix C Additional Evaluation Results
@@
Membership risks of NODEs, based on ODENet-64 architecture. Table 3 and Table 4 summarizes our additional results on NODE (ODENet-64) models and NSDEs with $k\!=\!0$, respectively. It is notable that ODENet-64 demonstrates comparable results to ODENet-16_32_64 but with a decrease in test accuracy of 1.64% and 2.41% on CIFAR-10 and CIFAR-100, respectively. It is 2.5-6 $\times$ less susceptible to LiRA attacks compared to ResNet-14. On CIFAR-100, ODENET exhibits superior defense capabilities against LiRA attacks than ODENET-16_32_64, demonstrating a lower TPR@0.1%FPR of 1.72%, thereby demonstrating enhanced security against membership inference attacks compared to ResNet-14. For NSDEs with $k\!=\!0$, despite being the non-privacy variant, it exhibits robust defense capabilities. NSDEs with $k\!=\!0$ maintains competitive test accuracy and generalization gaps on both CIFAR datasets. Notably, NSDEs $k\!=\!0$ showcases it strengths in defending against membership inference attacks, with TPR@0.1%FPR of 0.35% and 1.03% for CIFAR-10 and CIFAR-100, respectively, demonstrating a stronger defense compared to ResNet-14 and NODE models. This superior performance in TPR@0.1%FPR, coupled with robust accuracy metrics, underscores NSDE’s ability to enhance security against advanced threats.
Figure 8. Membership risks and overfitting in CIFAR-100.
@@
LiRA success vs. generalization gap in CIFAR-100. Figure 8 illustrates the relationship between membership risks measured by LiRA and the generalization gap for models trained on CIFAR-100. For ResNet-14, the TPR at 0.1% FPR ranges from 8%–14%, even as the generalization gap remains consistently around 40%. This suggests that overfitting contributes to membership inference risks. In contrast, NODEs (ODENet-16_32_64 and ODENet-64) show slightly lower generalization gaps, typically ranging from 5%–30%. Their TPR at 0.1% FPR values are also lower, fluctuating betweeen 0.5%–5.5%. This suggests that NODEs are less prone to overfitting, thereby reducing their risks to membership inference attacks.
@@
Figure and tables for additional results. Here, we present full experimental results to support the findings reported in the main body of the manuscript: Tables 59, where each table covers specific aspects. Table 5 demonstrates the results with varying number of blocks of models in Figure 2 and 6. Table 6 and 7 provide supporting data for Figure 2, illustrating the results of the impact of solvers and variants for NODEs. The detailed evaluation results of defenses, which provide an in-depth analysis corresponding to Figure 3, are presented in Table 8. A complete set of the ablation study regarding the hyperparameters characterizing NSDEs are listed in Table 9, where the subset of it is reported in Figure 6.
Table 4. Membership inference risks of SDENet $\sigma\!=\!0$ models. We evaluate SDENet $\sigma\!=\!0$ models against five existing attacks on CIFAR-10 and CIFAR-100. We compare their risks based on four different metrics used in the prior work.
@@
Acc. (C-10) Acc. (C-100) TPR @ 0.1% FPR TPR @ 1% FPR AUC Inference acc. Model Method Train Test Train Test C-10 C-100 C-10 C-100 C-10 C-100 C-10 C-100 SDENet $\sigma\!=\!0$ Yeom et al. (Yeom et al., 2018) 89.09% 83.93% 68.01% 52.49% 0.00% 0.09% 0.39% 1.12% 0.538 0.625 54.00% 60.52% Shokri et al. (Shokri et al., 2017) 0.07% 0.05% 0.68% 0.63% 0.468 0.449 50.02% 50.00% Song and Mittal (Song et al., 2021) 0.00% 0.14% 0.50% 1.40% 0.516 0.496 51.65% 51.13% Watson et al. (Watson et al., 2022) 0.00% 0.09% 0.44% 1.12% 0.538 0.625 54.00% 60.52% Carlini et al. (Carlini et al., 2022) 0.35% 1.03% 2.02% 5.36% 0.559 0.676 54.09% 62.26%
@@
Table 5. Comparing membership risks of NODEs and NSDEs. We vary the number of blocks in {1, 2, 3} and use LiRA.
Acc. (C-10) Acc. (C-100) TPR @ 0.1% FPR TPR @ 1% FPR AUC Inference acc. Model # of Blocks Train Test Train Test C-10 C-100 C-10 C-100 C-10 C-100 C-10 C-100 ODENet-16_32_64 1 94.30% 84.41% 82.05% 52.16% 1.01% 3.30% 4.22% 12.11% 0.616 0.782 57.65% 72.88% 2 96.39% 85.26% 93.4% 52.8% 1.57% 7.01% 5.95% 20.86% 0.638 0.848 58.99% 75.77% 3 96.82% 85.25% 94.71% 53.07% 1.66% 7.78% 6.17% 22.38% 0.641 0.852 58.86% 76.32% ODENet-64 1 94.14% 84.36% 64.54% 70.93% 0.86% 1.72% 4.17% 7.70% 0.639 0.741 59.57% 69.49% 2 97.07% 85.20% 80.74% 53.39% 1.25% 2.41% 5.43% 10.59% 0.656 0.788 60.78% 64.33% 3 97.74% 85.82% 87.26% 53.78% 1.49% 3.29% 6.05% 13.71% 0.669 0.815 60.89% 69.85% SDENet-16_32_64 1 81.89% 81.91% 56.54% 50.64% 0.14% 0.43% 1.24% 3.00% 0.525 0.612 51.89% 57.64% 2 82.41% 82.42% 57.2% 51.12% 0.14% 0.45% 1.25% 2.99% 0.528 0.619 52.23% 58.57% 3 74.44% 74.96% 55.73% 48.27% 0.08% 0.32% 0.90% 2.58% 0.511 0.604 50.00% 57.43%
@@
Table 6. Membership risks of NODEs trained with different solvers. We compare the LiRA success on them.
@@
Table 7. Membership risks of NODE variants. We compare the LiRA success on these models in CIFAR-10.
@@
Table 8. Effectiveness of defenses in CIFAR-10. We use LiRA. All defenses, except for SDENets, are applied to ResNet-14..
@@
Table 9. Full evaluation results of different NSDE configurations.

### v2

@@ B. Differential Privacy of Gaussian Mechanism
The following result is proved in [44]. Since it is crucial to the results in the body, a slightly different proof is provided here for completeness.
Proposition 1 (Theorem A.1 in [44]).
@@
$\Pr\!\left[Z>t-\tfrac{\varepsilon}{2t}\right]\leq\tfrac{1}{\sqrt{2\pi}(t-\varepsilon/2t)}\exp\!\left(-\tfrac{1}{2}(t-\varepsilon/2t)^{2}\right)<\delta/2$
@@
Note that, by enforcing $t\geq 1$ and $10<\varepsilon<$ 1, the first term on the left-hand side is nonnegative when $t-\frac{\varepsilon}{2t}\geq t-\frac{1}{2}\geq 1$. Therefore, it suffices to choose $t\geq 3/2$ to ensure positivity. On the other hand, the second term is increasing in $t$ when $t>\sqrt{\varepsilon/2}$ and decreasing in $\varepsilon$ when $\varepsilon<2t^{2}$, so it follows that we may ensure
$(t\!-\!\tfrac{\varepsilon}{2t})^{2}=t^{2}\!-\!\varepsilon\!+\!(\tfrac{\varepsilon}{2t})^{2}\geq t^{2}\!-\!1\!+\!1\cdot 3^{-2}=t^{2}\!-\!\tfrac{8}{9}>2\ln(\tfrac{1}{\delta}\sqrt{\tfrac{2}{\pi}}).$
@@
$t^{2}>2\ln\left(\frac{1}{\delta}\sqrt{\frac{2}{\pi}}\right)+\ln\left(e^{8/9}\right)=2\ln\left(\frac{1}{\delta}\right)+\ln\left(e^{8/9}\frac{2}{\pi}\right)$
@@
and since $(2/\pi)e^{8/9}<1.25^{2}$, this condition is satisfied when $t^{2}>2\ln\left(\frac{1.25}{\delta}\right)$. This shows that the privacy loss random variable is bounded above by $\varepsilon$ with probability $1-\delta$ on ${}q,~{}q^{\prime}$ whenever $\sigma\geq\sqrt{2\ln\left(1.25/\delta\right)}\allowbreak(\left|~{}v\right|/\varepsilon)$. To remove the dependence on ${}q,~{}q^{\prime}$, it suffices to choose $\sigma\geq\sqrt{2\ln\left(1.25/\delta\right)}(S_{~{}f}/\varepsilon)$. ∎
C. Additional Evaluation Results
@@
Membership risks of NODEs, based on ODENet-64 architecture. Table V summarizes our additional results on NODE (ODENet-64) models and NSDEs with $\sigma\!=\!0$, respectively. It is notable that ODENet-64 demonstrates comparable results to ODENet-16_32_64 but with a decrease in test accuracy of 1.64% and 2.41% on CIFAR-10 and CIFAR-100, respectively. It is 2.5-6 $\times$ less susceptible to LiRA attacks compared to ResNet-14. On CIFAR-100, ODENet exhibits superior defense capabilities against LiRA attacks than ODENet-16_32_64, demonstrating a lower TPR@0.1%FPR of 1.72%, thereby demonstrating enhanced security against membership inference attacks compared to ResNet-14. For NSDEs with $\sigma\!=\!0$, despite being the non-privacy variant, it exhibits robust defense capabilities. NSDEs with $\sigma\!=\!0$ maintains competitive test accuracy and generalization gaps on both CIFAR datasets. Notably, NSDEs $\sigma\!=\!0$ showcases it strengths in defending against membership inference attacks, with TPR@0.1%FPR of 0.35% and 1.03% for CIFAR-10 and CIFAR-100, respectively, demonstrating a stronger defense compared to ResNet-14 and NODE models. This superior performance in TPR@0.1%FPR, coupled with robust accuracy metrics, underscores NSDE’s ability to enhance security against advanced threats.
Figure 8: Membership risks and overfitting in CIFAR-100.
@@
Impact of generalization gap in attack success (CIFAR-100). Figure 8 illustrates the relationship between membership risks measured by LiRA and the generalization gap for models trained on CIFAR-100. For ResNet-14, the TPR at 0.1% FPR ranges from 8%–14%, even as the generalization gap remains consistently around 40%. This suggests that overfitting contributes to membership inference risks. In contrast, NODEs (ODENet-16_32_64 and ODENet-64) show slightly lower generalization gaps, typically ranging from 19.08%–30%. Their TPR at 0.1% FPR values are also lower, fluctuating betweeen 0.5%–5.5%. This suggests that NODEs are less prone to overfitting, thereby reducing their risks to membership inference attacks.
Membership risks of replace-then-finetune strategy. We analyze the computational resources required for training ResNet, NODEs, NSDEs and NSDEs with replace-then-finetune models. As shown in Table VI, training NSDE-based models from scratch is computationally expensive. Table VII demonstrates the results with membership inference risks of replace-then-finetune strategy. We evaluate NSDEs with replace-then-finetune models against six existing attacks on four datasets. Although only replacing the final blocks of a ResNet with SDENet, the replace-then-finetune models narrows the generation gap while achieving lower TPR at both 0.1% and 1% FPR across four dataset, indicating improved robustness compared to the ResNet models.
@@
Figure and tables for additional results. Here, we present full experimental results to support the findings reported in the main body of the manuscript: Tables VIIIXI, where each table covers specific aspects. Table VIII demonstrates the results with varying number of blocks of models in Figure 2 and 6. Table IX and X provide supporting data for Figure 2, illustrating the results of the impact of solvers and variants for NODEs. The detailed evaluation results of defenses, which provide an in-depth analysis corresponding to Figure 3, are presented in Table XI.
TABLE VI: Computational resources required for training on CIFAR-10 and CIFAR-100.
Model Time per epoch Memory usage ResNet-14 10s 941MB NODE (ODENet-16-32-64)_block1 23s 1305MB NSDE (SDENet $k\!=\!2$)_block1 30s 5051MB NSDE (replace-then-finetune) 15s 1575MB
TABLE VII: Membership risks of replace-and-finetune strategy ResNet+SDENet.
@@
TPR @ 0.1% FPR TPR @ 1% FPR AUC Inference acc. Method F-M C-10 C-100 T-I F-M C-10 C-100 T-I F-M C-10 C-100 T-I F-M C-10 C-100 T-I Yeom et al. [3] 0.01% 0.00% 0.07% 0.09% 0.06% 0.14% 1.10% 1.12% 0.558 0.555 0.646 0.638 55.72% 55.81% 64.41% 61.22% Shokri et al. [4] 0.06% 0.05% 0.05% 0.12% 0.65% 0.52% 0.56% 1.04% 0.477 0.450 0.433 0.502 50.01% 50.00% 50.00% 50.20% Song and Mittal [38] 0.00% 0.01% 0.13% 0.13% 0.24% 0.24% 1.48% 1.19% 0.468 0.530 0.496 0.500 51.05% 53.29% 51.15% 50.36% Watson et al. [34] 0.00% 0.00% 0.07% 0.11% 0.10% 0.15% 1.10% 1.17% 0.550 0.555 0.646 0.537 53.81% 55.81% 64.41% 52.83% Carlini et al. [6] 1.39% 0.45% 1.65% 0.21% 4.48% 2.84% 7.51% 1.57% 0.576 0.600 0.726 0.535 53.27% 56.78% 67.37% 52.42% Zarifzadeh et al. [35] 0.73% 1.41% 2.91% 0.26% 4.74% 5.68% 10.88% 1.89% 0.544 0.607 0.743 0.561 54.80% 57.97% 68.67% 54.54%
@@
TABLE VIII: Comparing membership risks of NODEs and NSDEs. We vary the number of blocks in {1, 2, 3} and use LiRA.
Acc. (C-10) Acc. (C-100) TPR @ 0.1% FPR TPR @ 1% FPR AUC Inference acc. Model # of Blocks Train Test Train Test C-10 C-100 C-10 C-100 C-10 C-100 C-10 C-100 ODENet-16_32_64 1 94.30% 84.41% 82.05% 52.16% 1.01% 3.30% 4.22% 12.11% 0.616 0.782 57.65% 72.88% 2 96.39% 85.26% 93.40% 52.8% 1.57% 7.01% 5.95% 20.86% 0.638 0.848 58.99% 75.77% 3 96.82% 85.25% 94.71% 53.07% 1.66% 7.78% 6.17% 22.38% 0.641 0.852 58.86% 76.32% ODENet-64 1 94.14% 84.36% 70.93% 51.85% 0.86% 1.72% 4.17% 7.70% 0.639 0.741 59.57% 69.49% 2 97.07% 85.20% 80.74% 53.39% 1.25% 2.41% 5.43% 10.59% 0.656 0.788 60.78% 64.33% 3 97.74% 85.82% 87.26% 53.78% 1.49% 3.29% 6.05% 13.71% 0.669 0.815 60.89% 69.85% SDENet-16_32_64 1 81.89% 81.91% 56.54% 50.64% 0.14% 0.43% 1.24% 3.00% 0.525 0.612 51.89% 57.64% 2 82.41% 82.42% 57.20% 51.12% 0.14% 0.45% 1.25% 2.99% 0.528 0.619 52.23% 58.57% 3 74.44% 74.96% 55.73% 48.27% 0.08% 0.32% 0.90% 2.58% 0.511 0.604 50.00% 57.43%
@@
TABLE IX: Membership risks of NODEs trained with different solvers. We compare the LiRA success on them.
@@
TABLE X: Membership risks of NODE variants. We compare the LiRA success on these models in CIFAR-10.
@@
TABLE XI: Effectiveness of defenses in CIFAR-10. We use LiRA. All defenses, except for SDENets, are applied to ResNet-14.
@@
Figure 9: Membership inference risks of ODENet-64 models. We evaluate ODENet-64 models with varying numbers of NODE blocks against the attack of LiRA in CIFAR-10.
Figure 10: Membership risks of NSDEs train with different step-size ($1/s$). (Top row, left to right) We illustrate the impact of step-size ($1/s$) with the level of stochasticity $k\!=\!0.5$, and varying the integration Time $T\!=\!0.5,0.25,0.125$, respectively. (Bottom row, left to right) We illustrate the impact of step-size ($1/s$) with the level of noise intensity $\sigma\!=\!2$, and varying the integration Time $T\!=\!0.5,0.25,0.125$, respectively.
Figure 11: Membership risks of NSDEs train with different time interval ($T$). We illustrate the impact of time ($T$) with the level of stochasticity $k\!=\!0.5$ (left two) and noise intensity $\sigma\!=\!2$ (right two), and varying $s\!=\!16,32$, respectively.
TABLE XII: Full evaluation results of different NSDE configurations.